# Towards Reliable Human Evaluations in Gesture Generation: Insights from a Community-Driven State-of-the-Art Benchmark

### v2

@@ A.1 Study participants
Crowdworkers who agree to take the test are first presented with an introduction screen presenting brief instructions about the task, with example screenshots of the response options. After that follow 25 test pages, each presenting a pair of video stimuli along with buttons and tickboxes for test takers to provide their responses. An example of the test-taker GUI is presented in Fig. 7; the exact questions depend on the type of the user study, following Fig. 6.
A.2 Speech-Segment Selection
To particpiate in the evaluation, model authors were required to submit generated motion for the entire BEAT2 English test set, in long form (one long motion for each take in the test set). Submitters whose system is not deterministic, i.e., it may generate different output motion for the same inputs, are required to generate motion for the test set five times with different random seeds and submit all five generations, to allow deeper analysis of the diversity of their generated motion.
However, following prior large-scale evaluations in speech [65] and gestures [43, 45, 44], not all of the submitted motion was evaluated in the user-studies. Instead, a number of short speech segments – defined as start and end times in the test data – were selected, and only motion corresponding to these segments of speech are considered for the user studies. Using this one set of speech segments as the basis for all user studies carried out on the BEAT2 dataset, we control for the effect of the speech on system behaviour (e.g., the same speakers are always represented in the same way in every evaluation).
The criteria for selecting speech segments for the user studies were as follows:
  • Each segment should correspond to one or more complete sentences.
  • Segment duration should be within the range of 7.0 to 12.0 seconds.
  • Segments should be disjunct (no overlap).
Only keeping segments where the recorded motion is free of major artefacts is important, since our objective is to compare new gesture-generation methods against the realism of actual human motion, not against the realism achieved by the specific 3D motion-extraction methodology. (The presence of artefacts in the ostensibly natural motion in Yoon et al. [87], Kucherenko et al. [45] may explain their counter-intuitive finding that synthetic motion was preferred over natural motion for one of the systems they evaluated.) Concretely, we eliminated (1) all segments containing flicking, which we define as instant, physically impossible changes in pose, and most instances of (2) mesh penetration/self-intersection.
Although a lot of the test dataset contains somewhat awkward finger poses due to the difficulty of tracking fingers, this was deemed less visually distracting and would be more drastic to exclude, so it was not considered grounds for exclusion. The only cases where mesh penetration were permitted were (2a) when the penetration visually resembled clothing or tissue giving way to light pressure, or (2b) where the penetration occurred due to the aforementioned poor finger tracking, as long as the fingers at worst were merely seen clipping into each other, and not passing through each other to the other side at any point. We decided to retain segments satisfying (2b) for the evaluations because these instances of mesh penetration are associated with poses having gestural importance, such as interwoven hands or finger against palm, which are important not to exclude.
A subselection of about 250 randomly chosen initial segments, drawn equally from all speakers, was manually whittled down to a a final set of 108 segments to be used in the evaluations on the BEAT2 dataset, only keeping segments whose motion had no major artefacts as described above. Motion-capture quality varied significantly across speakers, so we chose eight segments for speakers Wayne and Scott (which had the highest motion-capture quality) and four good segments for every other speaker in the test set. The chosen segments were between 7.7 and 12.0 seconds in duration (average 10.7 seconds) were selected. 108 segments is a significantly larger number of segments than used in previous large evaluations in the field, specifically the GENEA Challenges (which used 40–48 segments), yet still few enough that each stimulus will be assessed numerous times in the user studies, so as to allow future papers to perform stimulus-level statistical analyses of the data produced by our evaluation.
A.3 Attention Checks
Consistent with best practices in crowdsourced evaluations, we use attention checks to ensure that test-takers are paying attention to the task. These take the form of a message “[Attention check] Please choose ‘R’.”, with R being one of the five response options underneath the videos, chosen at random. Four attention checks are inserted into each user study, evenly spaced from the 20% until the 80% progress mark. Test-takers that fail any attention check are removed from the statistical analyses; those that fail more than one are rejected without pay. (Prolific’s policies do not permit rejecting test takers due to a single failed attention check.)
A.4 Statistical Analysis for Motion Realism
@@
Our statistical analysis is based on the pairwise preference data described above. We standardise the condition labels to canonical forms and convert the raw choices into triplets of the form (model_A, model_B, winner), as required by our scoring algorithm. A “clear” preference counts as two wins for the winner, whereas a “slight” preference only counts as one; ties count as half a win and half a loss for both models in the presentation.
@@
A.5 Statistical Analysis of Speech Appropriateness
For the statistical analysis, we use the basically same setup as described in Sec. A.4 for motion realism: clear preference responses count double compared to slight preferences, and ties (“They are equal”) count as half a win and half a loss. The only difference is that wins for the matched stimulus are assigned the value 1 and wins for the mismatched stimulus are assigned a 0. The resulting average appropriateness score (essentially a modified win rate) is then a number between zero and one.
The rest of the analysis is the same as for motion realism. We use the exact same test-taker-level bootstrap methodology to obtain confidence intervals, based on quantiles of the bootstrap distribution. To test if differences in appropriateness score between two conditions are statistically significantly different from zero, we likewise look at the differences between the appropriateness scores of any two systems in the bootstrap samples. We compute two-sided $p$-values regarding the score difference in the same way as for motion realism and apply Benjamini-Hochberg FDR correction [10] afterwards.
A.6 JUICE scores for motion realism
@@
A.7 JUICE scores for gesture-speech appropriateness
@@
A.8 Gesture-Motion Visualisation
Following the approach in Yoon et al. [87], Kucherenko et al. [45], our visualisation (Fig. 10) includes full-body motion with root-node translation and rotation. This captures the character’s positioning and stance with respect to the camera, leading to a more lifelike and expressive visualisation than methods that restrict animation to only the upper body [43, 89, 62, 21].
The SMPL-X mesh offers an anatomically accurate body shape and proportion that matches each speaker [64]. Although SMPL-X includes gendered meshes, we opted to use the gender-neutral mesh for all characters, which is further modified by the per-speaker body shape parameters. This decision was made because the gendered meshes often resulted in increased self-intersection when visualising motion from the BEAT2 test set. However, the gender-neutral mesh lacks features such as hair and clothing and does not replicate realistic skin tones or facial characteristics. The male- and female-presenting speakers are distinguished through different SMPL-X textures: male characters use the default white shirt texture, and female characters use the pink one, matched to the perceived gender of the voice. These textures also introduce variation in skin tone. To enhance realism, we added a hair prop and applied a displacement map to the clothing, improving its appearance and reducing the flatness seen in the SMPL-X mesh. Due to the absence of gaze information and inaccurate lipsync in BEAT2, we covered the face with a white mask. This avoids distracting visual artefacts that could negatively influence test-taker perception.
Furthermore, we excluded the feet from view, cutting the frame at approximately knee level. This was a deliberate choice motivated by known issues with foot-ground interactions, such as sliding and ground penetration, which are both common in synthetic motion and easily detectable by human observers. If shown, these artefacts would likely overshadow the gestural qualities under evaluation, as crowdsourced participants tend to focus on the most visually salient errors (see, for instance, the results in Sec. A.6). By omitting the feet, we help raters concentrate on the gestures themselves, aligning with the primary evaluation objective.
It is also worth noting that while foot-ground contact issues are primarily governed by straightforward laws of physics rules and can potentially be resolved through post-processing, upper-body gesturing is much less dictated by physical laws and is instead rooted in communicative and cultural conventions. This arguably makes gesturing a deeper and more challenging problem to solve in the long term, compared to issues of character interaction with the ground plane.
For the record, while the human evaluations are conducted based on cropped visualisations, the automatic metrics described in Appendix C operate on the full-body pose and motion data.
For consistency, all authors submit motion in SMPL-X format [64], and we render the video stimuli. This centralised rendering ensures uniform visual quality and prevents discrepancies that might otherwise arise from system-specific rendering pipelines.
Appendix B Details on Systems Evaluated
@@ B.4 RAG-Gesture [58]
The system is inherently trained on BEAT2 dataset and follows its input representation, therefore no adaptation to the trained model is made. Specifically it generates hand, body, face motion along with the translation of the character from a single model. As the framework follows the temporal VAE structure, it also performs long-form motion generation in chunks of 10-second time windows through autoregressive rollout (Sec. B.3). Consequently, retrieval algorithm is not used for the overlapping motion frames and RAG is performed for the newly generated motion. For evaluation, LLM-driven Gesture Type algorithm is used for RAG.
B.5 AMUSE [21]
@@
AMUSE was developed on the BEAT2 SMPL-X data, with the same dataset splits that we use. However, there are two differences between the submission format and the data processing of the original model that require adaptation. First, AMUSE puts emphasis on upper-body gesticulation rather than locomotion, therefore it discards the eight lower-body joints of the SMPL-X body. This was resolved by augmenting the model outputs with static lower-body joints. Second, AMUSE can only generate 10-second motion sequences, while the submission system normally expects a single, coherent motion sequence for each test-set file. As a workaround, the AMUSE submission contains 7–12 second motion clips, corresponding to the full set of speech segments described in Sec. A.2. Clips shorter than 10 seconds were generated by padding the audio input with silence, and discarding surplus motion frames from the output. Clips longer than 10 seconds were artificially created by blending two clips, $c_{1}$ and $c_{2}$, using spherical linear interpolation (SLERP), where $c_{1}$ is generated on the first 10 seconds of the segment, and $c_{2}$ is generated from the last two seconds of $c_{1}$ and the remaining portion of the segment.
B.6 HoloGest [20]
@@
Appendix C Experiments on Automatic Metrics
@@
Table 3 presents the results on the test set of the BEAT2 dataset. RAG-Gesture shows strong performance for distribution-based motion-quality metrics (FGD, $\text{FD}_{g}$, $\text{FD}_{k}$), as well as BA. HoloGest shows the best SRGR and $\text{DIV}_{pose}$. Semantic Gesticulator yields the best FGD and richest run-to-run variability ($\text{DIV}_{sample}$). Note that the FGD values are not directly comparable to those in Liu et al. [49] due to differing data sizes (see Chong and Forsyth [23]): we used all audio in the test set and all five random samples submitted for each system to obtain as many data points as possible for better distribution fitting.
@@
Inspired by the GENEA Challenges [45], we also conducted a quantitative analysis of the correlation between automatic metric scores and subjective human ratings. Given the limited number of systems, the correlation analysis serves as a reference and should not be interpreted as providing strong evidence or conclusive findings. Specifically, we computed the correlation between Elo ratings (representing human preferences) and each automatic metric using Kendall’s $\tau$ rank correlation [39]. For human ratings on motion realism, all motion quality metrics exhibited moderate negative correlations (between 0.4-0.4 and 0.6-0.6, consistent with the findings for FGD in Kucherenko et al. [45]), while SRGR – despite being more closely related to speech-gesture appropriateness – showed the highest positive correlation (0.73). For speech-motion appropriateness, BA demonstrated a moderate correlation (0.5), whereas SRGR showed virtually no correlation (0.14-0.14). However, none of these correlations were statistically significant ($p<0.05$).
Overall, our findings highlight that whilst automatic metrics can provide useful insights and facilitate early evaluation, they remain insufficient to replace human evaluation in gesture generation. The discrepancies – especially in speech–gesture appropriateness – underscore the limitations of current objective measures and the continued necessity of human evaluation.

### v4

@@ B.3.1 Details on Participant Recruitment
B.3.2 Attention Checks
Consistent with best practices in crowdsourced evaluations, our protocol includes attention checks to ensure that test-takers are paying attention to the task. These take the form of a message “[Attention check] Please choose ‘R’.”, with R being one of the five response options underneath the videos, chosen at random. Four attention checks are inserted into each user study, evenly spaced from the 20% until the 80% progress mark. Test-takers that fail any attention check are removed from the statistical analyses; those that fail more than one are rejected without pay. (Prolific’s policies do not permit rejecting test takers due to a single failed attention check.)
@@ B.3.2 Attention Checks
Appendix C Additional Details on Our Benchmarking
C.1 Statistical Analysis for Motion Realism
@@
Our statistical analysis is based on the pairwise preference data acquired from the evaluations. We standardise the condition labels to canonical forms and convert the raw choices into triplets of the form $(\text{model}_{A},\text{model}_{B},\text{winner})$, as required by our scoring algorithm. A “clear” preference counts as two wins for the winner, whereas a “slight” preference only counts as one; ties count as half a win and half a loss for both models in the presentation.
@@
C.2 Statistical Analysis of Speech Appropriateness
For the statistical analysis, we use the basically same setup as described in Sec. C.1 for motion realism: clear preference responses count double compared to slight preferences, and ties (“They are equal”) count as half a win and half a loss. The only difference is that wins for the matched stimulus are assigned the value 1 and wins for the mismatched stimulus are assigned a 0. The resulting average appropriateness score (essentially a modified win rate) is then a number between zero and one.
The rest of the analysis is the same as for motion realism. We use the exact same test-taker-level bootstrap methodology to obtain confidence intervals, based on quantiles of the bootstrap distribution.
C.3 JUICE Scores for Motion Realism
@@
C.4 JUICE Scores for Speech-Gesture Appropriateness
@@
Appendix D Details on Systems Evaluated
@@ D.4 RAG-Gesture [56]
The system is inherently trained on BEAT2 dataset and follows its input representation, therefore no adaptation to the trained model is made. Specifically it generates hand, body, face motion along with the translation of the character from a single model. As the framework follows the temporal VAE structure, it also performs long-form motion generation in chunks of 10-second time windows through autoregressive rollout (Sec. D.3). Consequently, retrieval algorithm is not used for the overlapping motion frames and RAG is performed for the newly generated motion. For evaluation, LLM-driven Gesture Type algorithm is used for RAG.
D.5 AMUSE [19]
@@
AMUSE was developed on the BEAT2 SMPL-X data, with the same dataset splits that we use. However, there are two differences between the submission format and the data processing of the original model that require adaptation. First, AMUSE puts emphasis on upper-body gesticulation rather than locomotion, therefore it discards the eight lower-body joints of the SMPL-X body. This was resolved by augmenting the model outputs with static lower-body joints. Second, AMUSE can only generate 10-second motion sequences, while the submission system normally expects a single, coherent motion sequence for each test-set file. As a workaround, the AMUSE submission contains 7–12 second motion clips, corresponding to the full set of speech segments described in Sec. B.1. Clips shorter than 10 seconds were generated by padding the audio input with silence, and discarding surplus motion frames from the output. Clips longer than 10 seconds were artificially created by blending two clips, $c_{1}$ and $c_{2}$, using spherical linear interpolation (SLERP), where $c_{1}$ is generated on the first 10 seconds of the segment, and $c_{2}$ is generated from the last two seconds of $c_{1}$ and the remaining portion of the segment.
D.6 HoloGest [18]
@@
Appendix E Experiments on Automatic Metrics
@@
Table 3 presents the results on the test set of the BEAT2 dataset. RAG-Gesture shows strong performance for distribution-based motion-quality metrics (FGD, $\text{FD}_{g}$, $\text{FD}_{k}$), as well as BA. HoloGest shows the best SRGR and $\text{DIV}_{\text{pose}}$. Semantic Gesticulator yields the best FGD and richest run-to-run variability ($\text{DIV}_{\text{sample}}$). Note that the FGD values are not directly comparable to those in Liu et al. [47] due to differing data sizes (see Chong and Forsyth [21]): we used all audio in the test set and all five random samples submitted for each system to obtain as many data points as possible for better distribution fitting.
@@
Inspired by the GENEA Challenges [44], we also conducted a quantitative analysis of the correlation between automatic metric scores and subjective human ratings. Given the limited number of systems, the correlation analysis serves as a reference and should not be interpreted as providing strong evidence or conclusive findings. Specifically, we computed the correlation between Elo ratings (representing human preferences) and each automatic metric using Kendall’s $\tau$ rank correlation [38]. For human ratings on motion realism, all motion quality metrics exhibited moderate negative correlations (between 0.4-0.4 and 0.6-0.6, consistent with the findings for FGD in Kucherenko et al. [44]), while SRGR – despite being more closely related to speech-gesture alignment – showed the highest positive correlation (0.73). For speech-motion alignment, BA demonstrated a moderate correlation (0.5), whereas SRGR showed virtually no correlation (0.14-0.14). However, none of these correlations were statistically significant ($p<0.05$).
Overall, our findings highlight that whilst automatic metrics can provide useful insights and facilitate early evaluation, they remain insufficient to replace human evaluation in gesture generation. The discrepancies – especially in speech–gesture alignment – underscore the limitations of current objective measures and the continued necessity of human evaluation.